A box of mass 15 kg is pushed with a force of 24 N (right) a frictional force of 18 N (left ) opposes the motion of the box what is the acceleration of the box?
1
Expert's answer
2019-10-17T11:10:59-0400

We have:


"ma=F-F_f"

"(15)a=24-18"

"a=0.4\frac{m}{s^2}"